OpenStudy (anonymous):

State the amplitude, period, and phase shift for the function f(x)=3sin(2 theta + pi)

OpenStudy (anonymous):

\[f(x)=3\sin(2\theta + \pi)\]

OpenStudy (anonymous):

AMPLITUDE IS 3 PERIOD IS PI PHASE IS PI/2